OpenStudy (anonymous):

Jeannie is 8 years older than her younger sister April. When added together, their ages equal 198. How old is April?

OpenStudy (anonymous):

someone please help me :( it seems really simple i just dont know how to write it into a equation

OpenStudy (anonymous):

Jeannie - 8 years =age of april So j - 8 = a Jeannie's age + Aprils age = 198 So j+a=198 So we have two equations 1) j+a=198 and 2) j-8=a So plug the 2nd into the first for a j+a=j+j-8=198 2j=206 j=103 so a=j-8 = 103-8 = 95 April is 95.